For which of the following reaction kp=ke?I2+H2=_2HI3N2+3H2=_2NH3PCL5=_PCL3+CL2

Grade:12th pass

1 Answers

Vikas TU
6 years ago
We have the Pk and Kc relation as:
Kp = Kc(RT)^(n)
where,
n = products stichiometric coeff. – reactants stichiom. coeff.
and for Kp = Kc n = 0
Therfore for I2+H2 = 2HI3
only n = 2 – (1 + 1) = 0
It has same kp and Kc.